# COMPASS: Cross-embOdiment Mobility Policy via ResiduAl RL and Skill Synthesis

### v1

@@ V-A1 Curriculum on Goal Distance
V-A2 Critic State
@@
In reinforcement learning, a key consideration is the proper design of the actor and critic networks. In our default setup, the critic network uses the same policy state as the actor, under the assumption that the policy state already encapsulates sufficient information to learn accurate value functions. While this approach often works well, it depends on implicit information, which may make learning the critic network more challenging. To investigate potential improvements, we explored different configurations for the critic’s input. In one variant, we provided a depth image explicitly to the value function by employing a ResNet18 for feature extraction, then concatenating the resulting features with velocity and route embeddings to form the critic state. As shown in Table II, this strategy enhanced performance in office scenarios—where tall obstacles exist—but inadvertently reduced performance in warehouse environments presented with more low-lying obstacles. These findings suggest that the additional signal from the depth image can boost performance under certain conditions, indicating a promising direction for further exploration.
V-A3 Training Environment
@@
In another ablation study, we evaluate the impact of the training environment. Our default setup uses a combined environment consisting of warehouse, office, and lab scenarios. To assess the effect of environmental diversity, we also conducted a separate training run using only the warehouse environment with H1 robot. The results, shown in Table II, indicate that restricting training to warehouse scenarios slightly enhances performance in warehouse scenes but reduces performance in the office and combined settings. These findings underscore the value of diverse training environments for improving generalization. Notably, even though the office setting was excluded from this specialized training, the model still performs reasonably well when tested there, demonstrating a degree of environment generalizablity.
V-B Policy Distillation
In our final ablation study, we want to examine the policy distillation process. The default setup trains the distilled policy to minimize KL divergence on an unfiltered dataset that may include failure cases. Our first goal is to assess whether capturing the full policy distribution is more beneficial than merely imitating the mean. To this end, we trained another model using only an MSE loss on the mean predictions. As shown in Table III, this approach led to slightly inferior performance, particularly on the H1 robot, suggesting that KL divergence—which preserves variance—provides more flexibility for managing training noise and architectural differences.
We also investigated the role of dataset quality by constructing a new dataset from the same scenarios but excluding all failure cases, hypothesizing that this might yield better overall performance. Contrary to our expectations, the results in Table III improved only marginally. We suspect that removing failure cases also eliminates corner cases that could otherwise offer valuable learning opportunities for the generalist policy.
TABLE III: Ablation study on policy distillation
Embodiment Koss / Dataset Warehouse Single Rack Warehouse Multi Rack Office Combined
@@ V-B Policy Distillation
VI Robot Deployment
Building on X-Mobility, which has already demonstrated zero-shot Sim2Real transfer on edge devices, COMPASS retains the original backbone and simply replaces the IL policy network with the distilled generalist policy. This ensures that the final model remains lightweight enough for real-time inference on common embedded platforms. To confirm its efficiency, Table IV shows the TensorRT inference benchmark results on AGX Jeton Orin.
TABLE IV: Inference Latency and GPU Memory Usage
Model P50(ms) P95(ms) GPU Memory(MB)
Distilled Generalist 29.3389 29.3767 422
VII Discussions
Our findings suggest several key takeaways and directions for future work:
@@ VII Discussions
  • Hierarchical Mobility Stack: We observe a decreased success rate in environments such as offices and multi-rack warehouses, primarily for tasks requiring long-horizon route planning. This underscores the importance of a hierarchical mobility stack with graph-based route planner to handle more complex mobility scenarios.
VIII Conclusion
We have introduced a novel workflow, COMPASS, for cross-embodiment mobility that leverages imitation learning, residual RL, and policy distillation. The initial IL step provides a strong baseline, while residual RL quickly adapts the model to each embodiment’s dynamics, and a final distillation step merges specialized policies into a single, generalist policy. Our experiments show that this approach scales effectively to various robot platforms and environments. Future investigations will focus on Sim2Real deployments, more sophisticated embodiment embeddings, and broader cross-embodiment settings into the same policy workflow.

### v3

@@ 5.1 Residual RL
5.1.1 Curriculum on Goal Distance
In our RL training setup, we currently do not use a curriculum. To investigate whether such a strategy could improve training efficiency, we introduced a curriculum based on goal distance. Rather than enforcing a minimum goal distance from the outset, we gradually increased the probability of applying this constraint as training progressed to enable smoother convergence.
However, the results in Table 3 indicate that the curriculum-based approach did not outperform the baseline without a curriculum. A likely explanation is that the IL base policy already provides a strong initialization, and reducing exposure to challenging scenarios with the curriculum may have limited the policy’s capacity to explore and succeed on more difficult tasks.
5.1.2 Critic State
@@
In our default setup, the critic network uses the same policy state as the actor, under the assumption that the policy state already encapsulates sufficient information to learn accurate value functions. While this approach often works well, it depends on implicit information, which may make learning the critic network more challenging. To investigate potential improvements, we explored different configurations for the critic’s input. In one variant, we provided a depth image explicitly to the value function by employing a ResNet18 for feature extraction, then concatenating the resulting features with velocity and route embeddings to form the critic state. As shown in Table 3, this strategy enhanced performance in office scenarios—where tall obstacles exist—but inadvertently reduced performance in warehouse environments presented with more low-lying obstacles. These findings suggest that the additional signal from the depth image can boost performance under certain conditions, indicating a promising direction for further exploration.
5.1.3 Training Environment
@@
In another ablation study, we evaluate the impact of the training environment. Our default setup uses a combined environment consisting of warehouse, office, and lab scenarios. To assess the effect of environmental diversity, we also conducted a separate training run using only the warehouse environment with H1 robot. The results, shown in Table 3, indicate that restricting training to warehouse scenarios slightly enhances performance in warehouse scenes but reduces performance in the office and combined settings. These findings underscore the value of diverse training environments for improving generalization. Notably, even though the office setting was excluded from this specialized training, the model still performs reasonably well when tested there, demonstrating a degree of environment generalizability.
Table 4: Ablation study on policy distillation
Emb. Koss / Dataset Warehouse Single Warehouse Multi Office Combined
@@
5.2 Policy Distillation
In our final ablation study, we want to examine the policy distillation process. The default setup trains the distilled policy to minimize KL divergence on an unfiltered dataset that may include failure cases. Our first goal is to assess whether capturing the full policy distribution is more beneficial than merely imitating the mean. To this end, we trained another model using only an MSE loss on the mean predictions. As shown in Table 4, this approach led to slightly inferior performance, particularly on the H1 robot, suggesting that KL divergence—which preserves variance—provides more flexibility for managing training noise and architectural differences.
We also investigated the role of dataset quality by constructing a new dataset from the same scenarios but excluding all failure cases, hypothesizing that this might yield better overall performance. Contrary to our expectations, the results in Table 4 improved only marginally. We suspect that removing failure cases also eliminates corner cases that could otherwise offer valuable learning opportunities for the generalist policy.
6 Applications
This section presents selected applications of the COMPASS workflow and its trained policy, demonstrating capabilities that extend beyond the COMPASS workflow itself.
6.1 Open Vocabulary Object Navigation
The COMPASS policy is trained for point-based navigation, where the navigation command specifies a target pose. Thanks to its flexible interface design, COMPASS can also support open-vocabulary object navigation by integrating it with an object localization model such as Locate3D [20] (see Fig. 8). Given a language prompt describing the target object, the object localization model predicts the corresponding 3D bounding box. Combined with user-defined parameters, such as relative heading and clearance distance, the target pose generator computes an appropriate goal pose for COMPASS to navigate toward. This setup enables a modular framework that extends COMPASS from point-based navigation to open-vocabulary object navigation.
Figure 8: Open vocabulary object navigation by integrating COMPASS and object localization.
6.2 Synthetic Datasets for VLA Fine Tuning
Given the high quality of the RL specialist policy, it can also be leveraged to generate large-scale synthetic datasets in Isaac Lab for training foundational Vision-Language-Action (VLA) models for navigation tasks that typically demand substantial amounts of diverse training data. As a demonstration, we post-trained the Gr00t N1.5 model [21] using the COMPASS distillation dataset generated by the G1 expert policy. In this setup, goal information and measured velocity were provided as state inputs, while ego-centric video frames served as the visual input, paired with a generic text prompt (“Robot Navigation Task”). The Gr00t model was then trained to predict the corresponding velocity commands.
Notably, despite being trained on only 600 trajectories and without any prior pre-training on navigation, the fine-tuned Gr00t model successfully acquired robust navigation capabilities. When evaluated on 640 randomly generated test cases in Isaac Lab, the post-trained policy achieved task success rates comparable to the original COMPASS policy in in-distribution warehouse environments, as shown in Table 5. Moreover, it significantly outperformed COMPASS in the hospital environment, which represents a fully out-of-distribution setting. The model also demonstrated seamless transfer to real-world environments without any additional fine-tuning, enabling smooth integration with the Gr00t manipulation policy for more advanced loco-manipulation tasks, as illustrated in Fig. 9.
Table 5: Success Rate of Gr00t N1.5 Post-Trained on the COMPASS Dataset
Model Warehouse Hospital
Gr00t N1.5 86.1% 77.6%
COMPASS 84.7% 45.6%
Figure 9: Zero-shot Sim2Real deployment of the post-trained Gr00t model for loco-manipluation tasks.
7 Future Works
We have introduced a novel workflow, COMPASS, for cross-embodiment mobility that leverages imitation learning, residual RL, and policy distillation. The initial IL step provides a strong baseline, while residual RL quickly adapts the model to each embodiment’s dynamics, and a final distillation step merges specialized policies into a single, generalist policy. Our experiments show that this approach scales effectively to various robot platforms and environments.
Our findings also suggest several key takeaways and directions for future work:
@@ 7 Future Works
  • Hierarchical Mobility Stack: We observe a decreased success rate in environments such as offices and multi-rack warehouses, primarily for tasks requiring long-horizon route planning. This underscores the importance of a hierarchical mobility stack with a graph-based route planner to handle more complex mobility scenarios.
8 Acknowledgment
We thank Di Zeng, Chirag Majithia, Sameer Chavan, Billy Okal, Dennis Da, Fernando Castaneda Garcia-Rozas, Peter Varvak, Manuel Chavez Orozco, Vishal Kulkarni, Qi Wang, Yuke Zhu, and Jim Fan for their assistance with system integration and robot testing and for providing insightful feedback on the workflow setup.